OpenStudy (anonymous):

just started on this unit help please:) For what values of b is f(x) = bx a decreasing function???

OpenStudy (anonymous):

A. Any real number B. b > 0 C. b > 1 D. 0 < b < 1

OpenStudy (anonymous):

b could be gradient negative is decreasing...im not sure...

OpenStudy (jdoe0001):

think about it this way b > 0 y = 5x x = 3, y = 15 y = 6x x = 4, y = 24 y = 7x x = 5, y = 35 ... now b < 0 y = -5x x = 3, y = -15 y = -6x x = 4, y = -24 y= -7x x = 5, y = -35 ... off those 2 lines, which one do you think is showing a "decreasing function graph"?

OpenStudy (anonymous):

I would say b<0 is decreasing

OpenStudy (anonymous):

but that isn't one of the choices:/

OpenStudy (anonymous):

confusion

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

now let's check if 0 < b < 1 y = 5/10x x = 3, y = 15/10 y = 6/10x x = 4, y = 24/10 y = 7/10x x = 5, y = 35/10 ... y = -5/10x x = 3, y = -15/10 y = -6/10x x = 4, y = -24/10 y = -7/10x x = 5, y = -35/10 ...

OpenStudy (jdoe0001):

so the function will seem to stick in the I and IV Quadrants only on the 1st quadrant is "ascending" on the 4th quadrant is "descending"

OpenStudy (jdoe0001):

... actually, my negative fractions are less than 0 :/

zepdrix (zepdrix):

Hmm ya your options don't seem to make sense.... Was the function in question suppose to be, \(\large f(x)=b^x\)

OpenStudy (jdoe0001):

lemme retry that one

OpenStudy (jdoe0001):

now let's check if 0 < b < 1 y = 5/10x x = 3, y = 15/10 y = 6/10x x = 4, y = 24/10 y = 7/10x x = 5, y = 35/10 ... is always in the 1st quadrant, and always ascending for that scenario

OpenStudy (jdoe0001):

pauli is it an exponential function as zepdrix pointed out?